A molecule of DNA (deoxyribonucleic acid) is 2.07 μm long. The ends of the molecule become singly ionized: negative on one end, positive on the other. The helical molecule acts like a spring and compresses 1.02% upon becoming charged. Determine the effective spring constant of the molecule.
Answer #1

Say, the force stays constant during compression.

The ends of the molecules become singly ionized meaning the charge on both ends is "e".

[e = 1.6x10^-19 C ]

F = kQ²/r²
F = (8.99x109)(1.6x10-19)²/(2.07x10-6
F = 5.371 x 10-17 N

F = Kx
x = 0.0102(2.07x10^-6)

5.371 x 10-17 = K(0.0102(2.07x10^-6) )
K = 2.543 x 10 -9N/m